OpenStudy (driftracer305):

the probability that a dessert contains chocolate is 74%. the probability that a dessert contains both chocolate and nuts is 24%. find the probability that a randomly chosen chocolate dessert contains nuts.

OpenStudy (driftracer305):

@ganeshie8

OpenStudy (driftracer305):

i am guessing 0.74*0.24 = 0.1776 = 17.8%......... is that right?

ganeshie8 (ganeshie8):

P(nuts | chocolate) = P(chocolate AND nuts) / P(chocolate)

OpenStudy (driftracer305):

so 0.24*0.74 / 0.74......?

ganeshie8 (ganeshie8):

nope, you're given : P(chocolate AND nuts) = 24% = 0.24 P(chocolate) = 74% = 0.74 just divide

ganeshie8 (ganeshie8):

P(nuts | chocolate) = P(chocolate AND nuts) / P(chocolate) = 0.24/0.74 = ?

OpenStudy (driftracer305):

ahh ok i see......... but when i made a tree diagram.... i thought that for nuts you would put it as 24%..........

OpenStudy (driftracer305):

so what if it asked for just chocolate without nuts? @ganeshie8

ganeshie8 (ganeshie8):

then we need more info i guess..

OpenStudy (driftracer305):

ohh ok..... thank you sir...... @ganeshie8 thx a lot......

ganeshie8 (ganeshie8):

yw!

ganeshie8 (ganeshie8):

I think we can calculate just chocolate without nuts ! contingency table makes it clear to see :